# Drawing strategies in Strong Ramsey games for 3-uniform hypergraphs

### v1

@@ Abstract.
The Strong Ramsey game $\mathcal{R}(B,G)$ is a two player game with players $P_{1}$ and $P_{2}$, where $B$ and $G$ are $k$-uniform hypergraphs for some $k\geq 2$. $G$ is always finite, while $B$ may be infinite. $P_{1}$ and $P_{2}$ alternately color uncolored edges $e\in B$ in their respective color and $P_{1}$ begins. Whoever completes a monochromatic copy of $G$ in their own color first, wins the game. If no one claims a monochromatic copy of $G$ in a finite number of moves, the game is declared a draw. In this paper, we give an infinite set of 3-uniform hypergraphs $\{G_{t}\}_{t\geq 3}$, such that $P_{2}$ has a drawing strategy in the Strong Ramsey game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G_{t})$. This improves a result by David, Hartarsky and Tiba [undefe].
@@ 2. Notation, Definitions and Basic Results
All (hyper)graphs considered in this paper are undirected and simple. Hence, we denote kk-(hyper-)edges as $x_{1}x_{2}\ldots x_{k}$ instead of $\{x_{1},x_{2},\ldots,x_{k}\}$ for clarity of the presentation. In the Strong Ramsey game $\mathcal{R}(B,G)$, we will call $B$ the board and $G$ the target graph. We refer to $P_{1}$ as she and $P_{2}$ as he and we will often say that a player claims, picks or takes an edge instead of colors it in their respective color. For brevity, we sometimes refer to an edge claimed by $P_{1}$ or $P_{2}$ as $P_{1}$-edge or $P_{2}$-edge, respectively. We use a similar notion, for graphs, which a player has claimed. We might also say that a player has or has built a graph $H$, meaning that the player has claimed all edges of a copy of $H\subseteq B$. Moreover, we say a vertex $x\in B$ is fresh, if no player has claimed an edge incident to $x$. For a given target graph $G$, we say that $P_{1}$ has a threat or $P_{1}$ threatens $P_{2}$, if $P_{1}$ has claimed a copy $H\subseteq B$ of GeG-e for an edge $e\in E(G)$, such that the corresponding $\hat{e}\in B$ is not yet claimed by either player, and similarly for $P_{2}$. We will call $H$ a threat graph in that case. In order to be concise, we will not give a formal definition of strategy here. Instead, we refer the reader to [undefa, Appendix C]. We define $k]\coloneqq\{1,\dots,k\}$.
@@
In the course of this paper, it will be convenient to use time stamps in order to clarify which stage of the game we are talking about. Therefore, for a fixed point in time $T$, we define $E_{T}(P_{1})$ as the set of edges and $e_{T}(P_{1})$ as the number of edges that $P_{1}$ has claimed at $T$. For vertices $v,w\in B$, we let $d_{T}^{P_{1}}(v)$ be the number of edges from $E_{T}(P_{1})$, which are incident to $v$. Moreover, we define $\Delta_{T}^{P_{1}}(B)\coloneqq\max_{v\in B}\{d_{T}^{P_{1}}(v)\}$.
@@ 3. $\mathcal{R}(K_{\aleph_{0}}^{(3)},\hat{K}_{2,4}^{(3)})$ is a draw
Let $a_{1}$ be a fresh vertex. $P_{2}$ claims $cya_{1}$ and $P_{1}$ must block $P_{2}$’s threat by taking $cxa_{1}$, because she had no threat. If claiming $cxa_{1}$ provides her with a threat graph $H$, we have $cxa_{1}\in E(H)$. Note that every edge of $H$ is incident to the center and one of the main vertices. Since $d^{P_{1}}(a_{1})=1$, it can neither be a center nor a main vertex of $H$. So $c$ and $x$ must be the center and one of the main vertices of $H$, but every threat graph with a vertex of degree 1 contains a $\hat{K}_{2,3}^{(3)}$, which was excluded by the assumptions $\mathit{(ii)}$ and $\mathit{(iii)}$.
$P_{2}$ continues by claiming $cya_{i}$ for fresh vertices $a_{i}$ as long as the game does not end or he can win by claiming $cxa_{i}$. Assume for a contradiction that $P_{1}$ won the game eventually. Then, she must have obtained a threat graph $H$ by blocking with $cxa_{i}$. Using the same argument as above, $P_{1}$ must have claimed a copy of $\hat{K}_{2,3}^{(3)}$ of type $\mathit{(ii)}$ or $\mathit{(iii)}$, which was forbidden. Hence, $P_{2}$ has a drawing strategy. ∎
@@
Let $x,y$ be the main vertices and $c$ the center of the $\hat{K}_{2,3}^{(3)}$, which $P_{2}$ has built and assume that either $\mathit{(ii)}$ or $\mathit{(iii)}$ of the Distraction Lemma (3.2) does not hold. In this case, $P_{1}$ has claimed at least seven edges, such that none of them is incident to $y$, since $cxy\in E(P_{2})$. $P_{1}$ can not have claimed another $\hat{K}_{2,3}^{(3)}$ with either main vertex or center $y$, since this requires at least four more edges on the $y$-board. Hence, $P_{1}$ has not claimed a copy of $\hat{K}_{2,3}^{(3)}$ with center $c$ or $y$ and main vertex $y$ or $c$, respectively. If we exchange $x$ and $y$ in the Distraction Lemma (3.2), we obtain that $P_{2}$ can force at least a draw. ∎
@@
In this case, $P_{2}$ will build a threat graph, whose center $c$ he will choose as follows. Let $c\in\{a,b\}$ be the vertex at $T_{1}$, such that $|F_{c}|\geq 2$. If both $|F_{a}|\geq 2$ and $|F_{b}|\geq 2$, we choose $c$, such that either $c\notin e^{*}$ or both $a,b\in e^{*}$. We let $x\in\{a,b\}$ be the other vertex, which will become a main vertex of the threat graph $P_{2}$ builds.
@@
$P_{2}$ claims $cv_{2}v_{5}$ up to relabeling vertices. This yields him a $\hat{K}_{2,3}^{(3)}$. We have either $\hat{e}=cv_{1}v_{5}$ or $\hat{e}\neq cv_{1}v_{5}$. In the latter case, $P_{1}$ must have taken $cv_{1}v_{5}$ in her next move after $T_{3}$, because he wins on the next move otherwise.
If $\hat{e}\neq cv_{1}v_{5}$, let $T_{4}$ be the point in time after $P_{1}$ took $cv_{1}v_{5}$. We check the conditions of the Distraction Lemma (3.2) at $T_{4}$. Suppose $P_{1}$ has claimed a copy of a threat graph $H$. Since she had no threat a move ago (at $T_{3}$), we must have $cv_{1}v_{5}\in E(H)$. Since the only possible centers for $H$ are $z_{1},z_{2}$ and $z^{*}$, by choice of $c,v_{1},v_{5}$, $z^{*}$ must be the center of $H$. This is not possible as the $z^{*}$-board contains at most a $P_{1}$-star centered at $z_{1}$ and two additional edges of $P_{1}$. For conditions $\mathit{(ii)}$ and $\mathit{(iii)}$, observe that if $z^{*}\neq x$, then $x$ is contained in at most three edges of $P_{1}$. Hence, it can neither be a main vertex nor a center of a $\hat{K}_{2,3}^{(3)}$. If $z^{*}=x$, then $c$ and $v_{5}$ satisfy conditions $\mathit{(ii)}$ and $\mathit{(iii)}$ as the only possible centers for a $\hat{K}_{2,3}^{(3)}$ claimed by $P_{1}$ are $z_{1},z_{2}$ and $z^{*}$.
If $\hat{e}=cv_{1}v_{5}$, $P_{1}$ can begin a second threat phase, if she picks $z_{1}z_{2}w$ or $z_{1}z^{*}w$ for a vertex $w$. $P_{2}$ blocks accordingly as before until $P_{2}$ selects an edge $e^{*}$, which does not contain both $z_{1}$ and $z_{2}$ or both $z_{1}$ and $z^{*}$. This results in the board state from the last paragraph at $T_{4}$ up to the number of times, that $P_{1}$ threatened $P_{2}$ by taking $z_{1}z_{2}w$ or $z_{1}z^{*}w$, respectively. Since $P_{2}$ had a drawing strategy in this position, this completes Case 1.3.
@@
If $z_{1}z_{2}z_{3}\in E(H)$, we obtain a board state, which can be reached via Case 1.2 or Case 1.3 as follows. Suppose $H$ was a copy of $G^{-}$, then $P_{1}$ could have reached the same position by claiming $G^{-}$ in his first eight moves. After $P_{2}$ blocks her threat as in Case 1.3, $P_{1}$ takes $cv_{1}v_{5}$ and $P_{2}$ takes $cv_{2}v_{5}$. Similarly for Case 1.2, if $P_{1}$ built a copy of $K_{2,4}^{(3)}$ within her first eight moves instead.
@@ 4. $\mathcal{R}(K_{\aleph_{0}}^{(3)},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$
We choose $x_{C}\in\{x_{1},x_{2}\}$ as follows. If $e^{P_{1}}_{T_{1}}(G_{t})=2t$ and $m_{2}=x_{1}$ or $m_{2}=x_{2}$, we choose $x_{C}=x_{2}$ or $x_{C}=x_{1}$, respectively and therefore guarantee, that $x_{C}$ is not a main vertex of $H$. In all other cases, we choose $x_{C}\in\{x_{1},x_{2}\}$, such that $d^{P_{1}}_{T_{1}}(x_{C})$ is minimal. $x_{C}$ will be the center of the $K^{(3)}_{2,t}(t-2)$, which $P_{2}$ will build in order to apply the Distraction Lemma (4.1). We will denote the other vertex in $\{x_{1},x_{2}\}$ with $x_{M}$ and it will become a main vertex of the $K^{(3)}_{2,t}(t-2)$. Note that, the only case in which we do not choose $d^{P_{1}}_{T_{1}}(x_{C})$ to be minimal is, if $P_{1}$ has claimed a (2t1)(2t-1)-star on the $a_{1}$-board with $x_{C}$ as a leaf together with $a_{1}x_{C}x_{M}$. $P_{2}$ now picks $x_{C}zy_{1}$ for a fresh vertex $z$. Let $T_{2}$ be the point in time after $P_{1}$ has made $2t+1$ moves.
@@
If $P_{1}$ had a threat graph $H$ at $T_{3}$, then t1t-1 of the $t$ edges specified in Figure 8 must be contained in $H$. Since $\{a_{1},a_{2},a_{3}\}\cap\{x_{C},z,y_{i}\}=\emptyset$ for all $i\in\{t+1,\dots,2t-1\}$, in fact all such edges $x_{C}zy_{i}$ must be contained in $H$. By choice of $z$, $x_{C}$ must be the center of $H$. Therefore, $d_{T_{3}}^{P_{1}}(x_{C})=3t-1$ (it cannot be larger, because $x_{C}\notin\{a_{1},a_{2},a_{3}\}$) and $d_{T_{1}}^{P_{1}}(x_{C})=2t-1$. By choice of $x_{C}$, we have $d_{T_{1}}^{P_{1}}(x_{M})=2t-1$, so $P_{1}$ must have claimed 2t12t-1 edges of the form $x_{C}x_{M}v$ for $v\notin\{y_{1},\dots,y_{2t-1}\}$ at $T_{1}$. Since $|V(G_{t})|=2t+2$, but the edges $P_{1}$ has claimed cover at least $3t+1$ vertices of $B$, it is clear that there cannot be a threat. In order to verify $\mathit{(ii)}$ and $\mathit{(iii)}$ of the Distraction Lemma (4.1), observe that $z$ cannot be a center of a $K_{2,t}^{(3)}$ (with main vertex $x_{C}$) claimed by $P_{1}$, because it was chosen fresh in move $2t$ by $P_{2}$.
@@
Figure 9. The $a_{1}$-board from the perspective of $P_{1}$ at $T_{2}$. The left blob contains the vertices $v_{1},\dots,v_{s}$, the middle contains $v_{s+1},\dots,v_{s+s^{\prime}}$ and the right contains $v_{s+s^{\prime}+1},\dots,v_{s+s^{\prime}+s^{\prime\prime}}$. The dashed edges are the ones in $A$, which are claimed by neither player yet.
@@
Assume the game is played up to $T_{4}$, the point in time when $P_{1}$ has either won or it is $P_{2}$’s move and he cannot play according to the algorithm. Let $E_{1}$ be the set of $P_{1}$-edges for which $P_{2}$ has claimed an edge on her next turn due to part $\mathit{(i)}$ of the algorithm and similarly for $E_{2},E_{3}$. Assume for a contradiction that $P_{1}$ won by building a copy $H$ of $G_{t}$ with center $c^{*}$ and main vertices $m_{1}^{*},m_{2}^{*}$, such that $d^{P_{1}}_{T_{4}}(m_{1}^{*})\geq 2t-1$. We begin by proving the following claim.
@@
Assume the contrary. By the calculations above, we obtain $|A|=2(t-s_{1})+1$. Due to part $\mathit{(i)}$ of the algorithm, $P_{1}$ can connect at most $t-s_{1}$ vertices from $v_{s_{1}+1},\dots,v_{s_{1}+s_{2}}$ to both $a_{2}$ and $a_{M}$. By $\mathit{(ii)}$, she can connect any other vertex to at most one of $a_{2},a_{M}$ on the $a_{1}$-board. So $P_{1}$ can obtain at most a $K_{2,t}$ and hence cannot even claim a $K_{2,t+1}^{(3)}$ with center $a_{1}$ and main vertices $a_{2}$ and $a_{M}$.
Note that at $T_{4}$ we have
$E_{1}\cup E_{2}\subseteq\{a_{1}a_{2}v^{\prime}\ |\ v^{\prime}\in V(B)\backslash\{a_{M}\}\}\cup\{a_{1}a_{M}v^{\prime\prime}\ |\ v^{\prime\prime}\in V(B)\backslash\{a_{2}\}\}.$
In addition, $|E_{3}|\leq t$, because $P_{2}$ has won by playing according to the algorithm otherwise.
Which vertex can $c^{*}$ possibly be? We know that $d_{T_{4}}^{P_{1}}(c^{*})\geq 3t$, hence the options for $c^{*}$ are $a_{1},a_{2}$ and $a_{M}$. Both the $a_{2}$-board and the $a_{M}$-board contain an arbitrarily large star centered at $a_{1}$ and potentially edges from $E_{3}$. But since $|E_{3}|\leq t$, this is not enough to build a copy of $G$, so $c^{*}=a_{1}$. Again by a degree argument, we have $m_{1}^{*}=a_{2}$ without loss of generality. ($m_{1}^{*}=a_{M}$ is almost the same.) Since $m_{2}^{*}\neq a_{M}$ by the claim above, $|\{e\in E_{1}\cup E_{2}\ |\ m_{2}^{*}\in e\}|\leq 2$ and $d_{T_{4}}^{P_{1}}(m_{2}^{*})\geq t+1$, we must have
(1) $\big|\{e\in E_{T_{4}}(P_{1})\cap E(H)\ |\ m_{2}^{*}\in e\}\backslash E_{3}\big|\in\{1,2\},$
so one of $a_{1}a_{2}m_{2}^{*},a_{1}a_{M}m_{2}^{*}$ must be contained in $E(H)$. Since $a_{1}a_{2}m_{2}^{*}\notin E_{T_{4}}(P_{1})\cap E(H)$, we obtain $a_{1}a_{M}m_{2}^{*}\in E_{T_{4}}(P_{1})\cap E(H)$. Moreover, all edges of $E_{3}$ must contain $a_{1}$ and $m_{2}^{*}$, because $|\{e\in E_{T_{4}}(P_{1})\cap E(H)\ |\ a_{1},m_{2}^{*}\in e\}|\leq t$, otherwise. In this case, $a_{1}a_{M}m_{2}^{*}\notin E(H)$, since $a_{1}a_{2}a_{M}\notin E_{T_{4}}(P_{1})$ contradicting Equation 1 and therefore our assumption.
So it must be $P_{2}$’s move and he cannot play according to the algorithm, due to $\mathit{(iii)}$. By the pigeonhole principle, $P_{2}$ has claimed the edges $x_{C}zy_{2},\dots,x_{C}zy_{t}$ without loss of generality and if he can not win at $T_{4}$, then $P_{1}$ must have claimed $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}$ in the meantime. Since $P_{2}$ has claimed a $K^{(3)}_{2,t}(t-2)$ and it is his move, we check the conditions of the Distraction Lemma (4.1).

### v2

@@ Abstract.
The Strong Ramsey game $\mathcal{R}(B,G)$ is a two player game with players $P_{1}$ and $P_{2}$, where $B$ and $G$ are $k$-uniform hypergraphs for some $k\geq 2$. $G$ is always finite, while $B$ may be infinite. $P_{1}$ and $P_{2}$ alternately color uncolored edges $e\in B$ in their respective color and $P_{1}$ begins. Whoever completes a monochromatic copy of $G$ in their own color first, wins the game. If no one claims a monochromatic copy of $G$ in a finite number of moves, the game is declared a draw. In this paper, we give an infinite family of 3-uniform hypergraphs $\{G_{t}\}_{t\geq 3}$, such that $P_{2}$ has a drawing strategy in the Strong Ramsey game $\mathcal{R}(K_{\aleph_{0}}^{(3)},G_{t})$. This improves a result by David, Hartarsky and Tiba [undefe].
@@ 2. Notation, Definitions and Basic Results
All (hyper)graphs considered in this paper are undirected and simple. Hence, we denote kk-(hyper-)edges as $x_{1}x_{2}\ldots x_{k}$ instead of $\{x_{1},x_{2},\ldots,x_{k}\}$ for clarity of the presentation. In the Strong Ramsey game $\mathcal{R}(B,G)$, we will call $B$ the board and $G$ the target graph. We refer to $P_{1}$ as she and $P_{2}$ as he and we will often say that a player claims, picks or takes an edge instead of colors it in their respective color. For brevity, we sometimes refer to an edge claimed by $P_{1}$ or $P_{2}$ as $P_{1}$-edge or $P_{2}$-edge, respectively. We use a similar notion for graphs which a player has claimed. We might also say that a player has or has built a graph $H$, meaning that the player has claimed all edges of a copy of $H\subseteq B$. Moreover, we say a vertex $x\in B$ is fresh, if no player has claimed an edge incident to $x$. For a given target graph $G$, we say that $P_{1}$ has a threat or $P_{1}$ threatens $P_{2}$, if $P_{1}$ has claimed a copy $H\subseteq B$ of GeG-e for an edge $e\in E(G)$, such that the corresponding $\hat{e}\in B$ is not yet claimed by either player, and similarly for $P_{2}$. We will call $H$ a threat graph in that case. In order to be concise, we will not give a formal definition of strategy here. Instead, we refer the reader to [undefa, Appendix C]. We define $k]\coloneqq\{1,\dots,k\}$.
@@
In the course of this paper, it will be convenient to use time stamps in order to clarify which stage of the game we are talking about. Therefore, for a fixed point in time $T$, we define $E_{T}(P_{1})$ as the set of edges and $e_{T}(P_{1})$ as the number of edges that $P_{1}$ has claimed at $T$. For vertices $v,w\in B$, we let $d_{T}^{P_{1}}(v)$ be the number of edges from $E_{T}(P_{1})$, which are incident to $v$.
@@ 3. $\mathcal{R}(K_{\aleph_{0}}^{(3)},\hat{K}_{2,4}^{(3)})$ is a draw
Let $a_{1}$ be a fresh vertex. $P_{2}$ claims $cya_{1}$ and $P_{1}$ must block $P_{2}$’s threat by taking $cxa_{1}$, because she had no threat. If claiming $cxa_{1}$ provides her with a threat graph $H$, we have $cxa_{1}\in E(H)$. Note that every edge of $H$ is incident to the center and one of the main vertices. Since $d^{P_{1}}(a_{1})=1$, it can neither be a center nor a main vertex of $H$. So $c$ and $x$ must be the center and one of the main vertices of $H$, but every threat graph with a vertex of degree 1 contains a $\hat{K}_{2,3}^{(3)}$, which was excluded by the assumptions (ii) and (iii).
$P_{2}$ continues by claiming $cya_{i}$ for fresh vertices $a_{i}$ as long as the game does not end and he cannot win by claiming $cxa_{i}$. Assume for a contradiction that $P_{1}$ won the game eventually. Then, she must have obtained a threat graph $H$ by blocking with $cxa_{i}$. Using the same argument as above, $P_{1}$ must have claimed a copy of $\hat{K}_{2,3}^{(3)}$ of type (ii) or (iii), which was forbidden. Hence, $P_{2}$ has a drawing strategy. ∎
@@
Let $x,y$ be the main vertices and $c$ the center of the $\hat{K}_{2,3}^{(3)}$, which $P_{2}$ has built and assume that either (ii) or (iii) of the Distraction Lemma (3.2) does not hold. In this case, $P_{1}$ has claimed at least seven edges not incident to $y$, since $cxy\in E(P_{2})$. $P_{1}$ cannot have claimed another $\hat{K}_{2,3}^{(3)}$ with either main vertex or center $y$, since this requires at least four more edges on the $y$-board. Hence, $P_{1}$ has not claimed a copy of $\hat{K}_{2,3}^{(3)}$ with center $c$ or $y$ and main vertex $y$ or $c$, respectively. If we exchange $x$ and $y$ in the Distraction Lemma (3.2), we obtain that $P_{2}$ can force at least a draw. ∎
@@
In this case, $P_{2}$ will build a threat graph; its center $c$ will be chosen as follows. Let $c\in\{a,b\}$ be the vertex at $T_{1}$, such that $|F_{c}|\geq 2$. If both $|F_{a}|\geq 2$ and $|F_{b}|\geq 2$, we choose $c$, such that either $c\notin e^{*}$ or both $a,b\in e^{*}$. We let $x\in\{a,b\}$ be the other vertex, which will become a main vertex of the threat graph $P_{2}$ builds.
@@
$P_{2}$ claims $cv_{2}v_{5}$ up to relabeling vertices. This yields him a $\hat{K}_{2,3}^{(3)}$ with main vertices $x$ and $v_{5}$. We have either $\hat{e}=cv_{1}v_{5}$ or $\hat{e}\neq cv_{1}v_{5}$. In the latter case, $P_{1}$ must have taken $cv_{1}v_{5}$ in her next move after $T_{3}$, because he wins on the next move otherwise.
If $\hat{e}\neq cv_{1}v_{5}$, let $T_{4}$ be the point in time after $P_{1}$ took $cv_{1}v_{5}$. We check the conditions of the Distraction Lemma (3.2) at $T_{4}$. Suppose $P_{1}$ has claimed a copy of a threat graph $H$. Since she had no threat a move ago (at $T_{3}$), we must have $cv_{1}v_{5}\in E(H)$. Since the only possible centers for $H$ are $z_{1},z_{2}$ and $z^{*}$, by choice of $c,v_{1},v_{5}$, $z^{*}$ must be the center of $H$. This is not possible as the $z^{*}$-board contains at most a $P_{1}$-star centered at $z_{1}$ and two additional edges of $P_{1}$. For conditions (ii) and (iii), observe that if $z^{*}\neq x$, then $x$ is contained in at most three edges of $P_{1}$. Hence, it can neither be a main vertex nor a center of a $\hat{K}_{2,3}^{(3)}$. If $z^{*}=x$, then $c$ and $v_{5}$ satisfy conditions (ii) and (iii) as the only possible centers for a $\hat{K}_{2,3}^{(3)}$ claimed by $P_{1}$ are $z_{1},z_{2}$ and $z^{*}$.
If $\hat{e}=cv_{1}v_{5}$, $P_{1}$ can begin a second threat phase, if she picks $z_{1}z_{2}w$ or $z_{1}z^{*}w$ for a vertex $w$. $P_{2}$ blocks accordingly as before until $P_{1}$ selects an edge, which does not contain both $z_{1}$ and $z_{2}$ or both $z_{1}$ and $z^{*}$. This results in the board state from the last paragraph at $T_{4}$ up to the number of times, that $P_{1}$ threatened $P_{2}$ by taking $z_{1}z_{2}w$ or $z_{1}z^{*}w$, respectively. Since $P_{2}$ had a drawing strategy in this position, this completes Case 1.3.
@@
If $z_{1}z_{2}z_{3}\in E(H)$, we obtain a board state, which can be reached via Case 1.2 or Case 1.3 as follows. Suppose $H$ was a copy of $G^{-}$, then $P_{1}$ could have reached the same position by claiming $G^{-}$ in her first eight moves. After $P_{2}$ blocks her threat as in Case 1.3, $P_{1}$ takes $cv_{1}v_{5}$ and $P_{2}$ takes $cv_{2}v_{5}$. Similarly for Case 1.2, if $P_{1}$ built a copy of $K_{2,4}^{(3)}$ within her first eight moves instead.
@@ 4. $\mathcal{R}(K_{\aleph_{0}}^{(3)},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$
We choose $x_{C}\in\{x_{1},x_{2}\}$ as follows. If $e^{P_{1}}_{T_{1}}(G_{t})=2t$ and $m_{2}=x_{1}$ or $m_{2}=x_{2}$, we choose $x_{C}=x_{2}$ or $x_{C}=x_{1}$, respectively and therefore guarantee that $x_{C}$ is not a main vertex of $H$. In all other cases, we choose $x_{C}\in\{x_{1},x_{2}\}$, such that $d^{P_{1}}_{T_{1}}(x_{C})$ is minimal. $x_{C}$ will be the center of the $K^{(3)}_{2,t}(t-2)$, which $P_{2}$ will build in order to apply the Distraction Lemma (4.1). We will denote the other vertex in $\{x_{1},x_{2}\}$ with $x_{M}$ and it will become a main vertex of the $K^{(3)}_{2,t}(t-2)$. Note that, the only case in which we do not choose $d^{P_{1}}_{T_{1}}(x_{C})$ to be minimal is, if $P_{1}$ has claimed a (2t1)(2t-1)-star on the $a_{1}$-board with $x_{C}$ as a leaf together with $a_{1}x_{C}x_{M}$. $P_{2}$ now picks $x_{C}zy_{1}$ for a fresh vertex $z$. Let $T_{2}$ be the point in time after $P_{1}$ has made $2t+1$ moves.
@@
If $P_{1}$ had a threat graph $H$ at $T_{3}$, then t1t-1 of the $t$ edges specified in Figure 8 must be contained in $H$. Since $\{a_{1},a_{2},a_{3}\}\cap\{x_{C},z,y_{i}\}=\emptyset$ for all $i\in\{t+1,\dots,2t-1\}$, in fact all such edges $x_{C}zy_{i}$ must be contained in $H$. By choice of $z$, $x_{C}$ must be the center of $H$. Therefore, $d_{T_{3}}^{P_{1}}(x_{C})=3t-1$ (it cannot be larger, because $x_{C}\notin\{a_{1},a_{2},a_{3}\}$) and $d_{T_{1}}^{P_{1}}(x_{C})=2t-1$. By choice of $x_{C}$, we have $d_{T_{1}}^{P_{1}}(x_{M})=2t-1$, so $P_{1}$ must have claimed 2t12t-1 edges of the form $x_{C}x_{M}v$ for $v\notin\{y_{1},\dots,y_{2t-1}\}$ at $T_{1}$. Since $|V(G_{t})|=2t+2$, but the edges $P_{1}$ has claimed cover at least $3t+1$ vertices of $B$, it is clear that there cannot be a threat. In order to verify (ii) and (iii) of the Distraction Lemma (4.1), observe that $z$ cannot be a center of a $K_{2,t}^{(3)}$ (with main vertex $x_{C}$) claimed by $P_{1}$, because it was chosen fresh in move $2t$ by $P_{2}$.
@@
Figure 9. The $a_{1}$-board from the perspective of $P_{1}$ at $T_{2}$. The left blob contains the vertices $v_{1},\dots,v_{s_{1}}$, the middle contains $v_{s_{1}+1},\dots,v_{s_{1}+s^{\prime}}$ and the right contains $v_{s_{1}+s^{\prime}+1},\dots,v_{s_{1}+s^{\prime}+s^{\prime\prime}}$. The dashed edges are the ones in $A$, which are claimed by neither player yet.
@@
Assume the game is played up to $T_{4}$, the point in time when $P_{1}$ has either won or it is $P_{2}$’s move and he cannot play according to the algorithm. In the latter case, all edges $x_{C}zy_{i}$ with $i\in[2t-1]$ must have been claimed at $T_{4}$. Let
[…]
and $\EE_{3}=E_{T_{4}}(P_{1})\backslash E^{\prime}$. Note that $E_{T_{2}}(P_{1})\subseteq E^{\prime}$. Moreover, after $P_{1}$ has claimed an edge $e\in E_{3}$, $P_{2}$ reacted according to (iii) in his next move, if possible. Hence, $|E_{3}|\leq t$ or $P_{2}$ has won by claiming $t+1$ edges of the form $x_{C}zy_{i}$ with $i\in[2t-1]$.
Assume for a contradiction that $P_{1}$ won by building a copy $H$ of $G_{t}$ with center $c^{*}$ and main vertices $m_{1}^{*},m_{2}^{*}$, such that $d^{P_{1}}_{T_{4}}(m_{1}^{*})\geq 2t-1$. We prove the following claim.
@@
Assume the contrary. By the calculations above, we obtain $|A|=2(t-s_{1})+1$. Due to part (i) of the algorithm, $P_{1}$ can connect at most $t-s_{1}$ vertices from $v_{s_{1}+1},\dots,v_{s_{1}+s_{2}}$ to both $a_{2}$ and $a_{M}$. By (ii), she can connect any other vertex to at most one of $a_{2},a_{M}$ on the $a_{1}$-board. So $P_{1}$ can obtain at most a $K_{2,t}$ on the $a_{1}$-board and hence cannot even claim a $K_{2,t+1}^{(3)}$ with center $a_{1}$ and main vertices $a_{2}$ and $a_{M}$.
Which vertex can $c^{*}$ possibly be? We know that $d_{T_{4}}^{P_{1}}(c^{*})\geq 3t$, hence the options for $c^{*}$ are $a_{1},a_{2}$ and $a_{M}$. Both the $a_{2}$-board and the $a_{M}$-board contain an arbitrarily large star centered at $a_{1}$ and potentially edges from $E_{3}$. But since $|E_{3}|\leq t$, this is not enough to build a copy of $G_{t}$, so $c^{*}=a_{1}$. Again by a degree argument, we have $m_{1}^{*}=a_{2}$ without loss of generality. ($m_{1}^{*}=a_{M}$ is almost the same.) Since $m_{2}^{*}\neq a_{M}$ by the claim above, we have $|\{e\in E^{\prime}\cap E(H)\ |\ m_{2}^{*}\in e\}|\leq 2$. Due to $|E_{3}|\leq t$ and $d_{T_{4}}^{P_{1}}(m_{2}^{*})\geq t+1$, we obtain that $|\{e\in E^{\prime}\cap E(H)\ |\ m_{2}^{*}\in e\}|\in\{1,2\}$, so one of $a_{1}a_{2}m_{2}^{*},a_{1}a_{M}m_{2}^{*}$ must be contained in $E(H)$. Since $c^{*}m_{1}^{*}m_{2}^{*}\notin E(H)$, we obtain $a_{1}a_{M}m_{2}^{*}\in E_{T_{4}}(P_{1})\cap E(H)$. Moreover, all edges of $E_{3}$ must contain $a_{1}$ and $m_{2}^{*}$, because $|\{e\in E_{T_{4}}(P_{1})\cap E(H)\ |\ a_{1},m_{2}^{*}\in e\}|\leq t$, otherwise. In this case, $a_{1}a_{M}m_{2}^{*}\notin E(H)$, since $a_{1}a_{2}a_{M}\notin E_{T_{4}}(P_{1})$ contradicting our assumption.
So it must be $P_{2}$’s move and he cannot play according to the algorithm, since all edges $x_{C}zy_{i}$ with $i\in[2t-1]$ have been claimed. By the pigeonhole principle, $P_{2}$ has claimed the edges $x_{C}zy_{2},\dots,x_{C}zy_{t}$ without loss of generality and if he cannot win at $T_{4}$, then $P_{1}$ must have claimed $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}$ in the meantime. Since $P_{2}$ has claimed a $K^{(3)}_{2,t}(t-2)$ and it is his move, we check the conditions of the Distraction Lemma (4.1).